OpenStudy (anonymous):

What is the energy of a photon of light having a wavelength of 4.5 x 10-7 meters? (Planck’s constant is 6.626 x 10-34 joule seconds; the speed of light is 2.998 x 108 m/s) 2.9817 x 10-40 joules 4.4143 x 10-19 joules 6.6745 x 10-26 joules second 9.9456 x 10-49 joules 19.864 x 10-26 joules/second

Vocaloid (vocaloid):

E = energy of photon = hc/lambda, where h is plank's constant, c is the speed of light, and lambda is the wavelength make sure all units are consistent

OpenStudy (michele_laino):

energy E carried by a photon is given by the subsequent formula: \[\Large E = h\nu \] where \nu is the frequency of our photon, namely \[\Large \nu = \frac{c}{\lambda }\]

OpenStudy (michele_laino):

c is the light speed

OpenStudy (michele_laino):

\lambda is the wavelength

Vocaloid (vocaloid):

@briannaboswell: energy of photon = E c = speed of light = 2.998*10^8 lambda = wavelength = 4.5 x 10-7 m h = plank's constant = 6.626 x 10-34 E = h*c/(lambda) all you have to do is plug in the values of h, c, and lambda...

OpenStudy (michele_laino):

so, substituting, we can write: \[\Large E = \frac{{hc}}{\lambda } = \frac{{6.62 \times {{10}^{ - 34}} \times 3 \times {{10}^8}}}{{4.5 \times {{10}^{ - 7}}}} = ...Joules\]